# Every Wrinkle Carries A Memory: An Integro-differential Bootstrap for Features in Cosmological Correlators

### v1

@@ 3.2 Monochromatic masses
around slow-roll backgrounds. Generically, similar oscillations would also appear in the cubic couplings, but to keep the problem’s complexity under control, we take the vertices to be time-independent, setting $\Omega_{L}=\Omega_{R}=0$. Nevertheless, monochromatic vertices could be incorporated into our proposed formulation subject to minor modifications.
@@
This object is proportional to the cubic wavefunction coefficient, $\psi_{3}=f(k_{12},s)/\sigma_{-}(s,\eta_{\rm end})$ and satisfies the same IDE as (3.18). With a slight misuse of terminology, we will refer to $f$ as the three-point function/building block. [^13] The precise relation to the three-point function is $\langle\phi(\bm{k}_{1})\phi(\bm{k}_{2})\sigma(-\bm{s},\eta_{0})\rangle^{\prime}=\frac{\eta_{\rm end}^{2}}{2k_{1}2k_{2}}\text{Re}\{f(k_{12},s)\sigma_{-}(s,\eta_{0})\}$.
@@ 3.3 Symmetries, microcausality and analyticity
The transformation rule for e.g. $F_{++}$ can be made manifest by writing [^14] Of course, the full four-point $F(k_{12},k_{34},s)$ admits the same decomposition. However, it is useful to decompose each Schwinger-Keldysh component separately for solving the integro-differential equations. {keyeqn}
@@
Inserting the above mode expansions into the respective IDEs does not yield any major simplification, other than producing a recursive set of IDEs for $F_{l}$ and $f_{l}$. Despite this recursive nature, the decomposition becomes especially convenient in perturbation theory, where—at a fixed order in $g^{2}$—only a finite number of harmonics need to be retained. To make this simplification more manifest, one can associate the ${\cal O}(g^{2l})$ contribution with a perturbative graph, as in Figure 3, consisting of $l$ mass insertions of the form $\Delta m^{2}(t)\,\sigma^{2}$. This structure implies the following scaling behaviour:
$F_{l}(u)\sim{\cal O}(g^{2|l|})\,,\qquad f_{l}(u)\sim{\cal O}(g^{2|l|})\,.$ (3.28)
This ensures, in particular, that only three harmonics ($l=0,\pm 1$) are needed to describe the exchange diagram $F_{++}$, and its three-point building block $f$, at linear order in $g^{2}$.
@@
where we have assumed $k_{12},k_{34}$ and $s$ all to be positive. Quite remarkably, the above factorisation property is non-perturbative in $g^{2}$ because it relies on the commutation of the heavy field operators outside the lightcone. All such commutators vanish independently of the mass time-dependence and the mode function, the detailed knowledge of which was not necessary to derive (3.33). However, our derivation relied on one critical assumption: the time integration itself does not introduce additional singularities in $\bm{s}$. This assumption is valid within a sufficiently small radius around $\bm{s}=0$, where the bulk-bulk propagator $G_{++}$ can be expanded into a power series in $s^{2}$, typically including both integer and fractional powers of $s^{2}$. Thanks to the $i\epsilon$ prescription, the corresponding Schwinger-Keldysh integral converges for $k_{12},k_{34}>0$—without introducing additional singularities near $\bm{s}=0$—even after substituting the propagator with this Taylor series. So non-analyticities in the Wightman propagator (at fixed $\eta,\eta^{\prime}$) are the only possible singularities for the integrated result $F_{++}$ around $\bm{s}=0$. We note in passing that (3.33) effectively generalises the cutting rule for non-local cosmological collider signals [69, 19, 71] to scale-breaking cases. Finally, let us highlight that, away from the origin, additional partial energy singularities emerge from time integration, as will be discussed shortly.
For computing the bispectrum and the power spectrum in our setup, it is sufficient to solve the seed exchange diagram $F_{++}$ only for soft configurations with $\bm{k}_{4}=0$ (hence $k_{34}=k_{3}$ and $s=k_{3}$). So it is convenient to have a factorisation theorem around $\bm{k}_{3}=0$ directly for $F(k_{12},k_{3},k_{3})$. With a small modification, $F(k_{12},k_{3},k_{3})$ turns out to factorise in the same way as (3.33): {keyeqn}
@@
where the second contribution is now analytic in $k_{3}$, not necessarily in $k_{3}^{2}$, as (3.33) would have naively suggested. Indeed, since this non-factorisable contribution originates from replacing the Feynman propagator $G_{++}$ with the retarded one $G_{R}$ in $F_{++}$, it inherits odd powers of $k_{3}$ from the right-vertex bulk-boundary propagator($\propto e^{ik_{34}\eta^{\prime}}=\,e^{ik_{3}\,\eta^{\prime}}$). In more detail, the time ordering $\eta^{\prime}>\eta$ enforced by the retarded propagator $G_{R}$ guarantees the convergence of the integral over $\eta^{\prime}$, even after substituting $e^{ik_{3}\,\eta^{\prime}}$ with its Taylor series within the Schwinger-Keldysh integrand. From this substitution, additional odd powers of $k_{3}=\sqrt{k_{3}^{2}}$ appear, multiplying even powers of $k_{3}$ already contained in $G_{R}$. This explains the difference between (3.33) and (3.34). Fortunately, odd powers of $k_{3}$ are the only type of non-analyticities that taking the soft limit $\bm{k}_{4}\to 0$ in $F_{++}$ can ever introduce; fractional powers of $k_{3}$ should still factorise in $F_{++}$ according to (3.34). This factorisation will serve as a powerful bootstrap input for solving the IDEs.
@@ 4.1 Three-point contact diagram
and [^17] Notice that $\hat{\Delta}_{u}$ here differs from its conventional definition in (1.1) by the inclusion of the mass term.
@@
The scale-invariant sector $f_{0}$, which is governed by (4.3), receives no contribution [^18] Note that, despite breaking the dilatation symmetry, mass oscillations starting at order $g^{4}$ do contribute to the scale-invariant part of the three-point function because with an even number of mass insertions in the perturbative expansion of $f$, the comoving scale $s$ can partially cancel out between the oscillatory vertices. This will lead to a finite contribution to the scale-invariant component $f_{0}(u)$ at each even order in $g^{2}$. at this order from mass oscillations. Therefore, it is equal to the corresponding three-point function in de Sitter (see e.g. [46, 1]),
@@
which is carried over [^20] In fact, both coefficients in (4.23) are exponentially suppressed but for different reasons: the $u^{1/2+i\mu}$ term is small because it is proportional to $\beta_{0}$, while the integration over the late time oscillations of the heavy field suppresses the coefficient of the $u^{1/2-i\mu}$ term. to the prefactor of the soft term $u^{1/2+i\mu}$ in (4.23) by integration over time. As a result,
@@
is also exponentially small.
@@
Let us focus on the mass range $H\ll m_{0}\lesssim\omega$. In this parametric regime, the Bogoliubov coefficient $\Delta\beta$ need not be exponentially small since particles even heavier than $H$ can still be produced by drawing energy from mass oscillations (equivalently, from the inflaton’s kinetic term). Indeed, the explicit formula for $\chi^{-}_{1}$ shows a power-law enhancement in $\Delta\beta_{-}$:
@@ 4.2 Three-point exchange diagram
    The particular solution (4.45) represent two qualitatively different contributions to the exchange diagram, $F_{++}=x_{0}^{\pm i\omega}F_{\pm}^{\text{ansatz}}$, as a function of the intermediate momentum $k_{3}$. The first line’s contribution is analytic around $k_{3}=0$ since factors of $k_{3}^{\pm i\omega}$ are precisely cancelled between the prefactor $x_{0}^{\pm i\omega}$ and $F^{\text{ansatz}}_{\pm 1}$. However, the second line in (4.45) induces a branch point at $k_{3}=0$ in the final four-point function. We shall soon come to the physical interpretation of each analytic behaviour below.
@@
    in which $_{p}\tilde{F}_{q}$ is the regularised hypergeometric function defined in (1.5). Let us stress that while it was sufficient to take the leading soft limit $k_{3}\to 0$ to fix the free coefficients $\xi_{1,2}$, the factorisation in (3.34) holds to all orders in $k_{3}$. We shall elaborate on this point below. Additional intermediate steps leading to the above results are spelled out in Appendix B.
@@
while $\mathcal{A}^{\pm}_{n>0}$ are of order $1/\mu^{4}$ or higher. By plugging the asymptotic form of $\mathcal{A}^{\pm}_{0}$ into (4.45) and neglecting the remaining terms, we recover the leading order EFT term $F^{\text{EFT}}_{\pm 1}$. Similarly, higher order EFT contributions could be isolated in (4.45) by keeping more powers of $1/\mu^{2}$ in the prefactors $\mathcal{A}^{\pm}_{n}$, and by expanding $_{3}F_{2}$ near U=0U=0.212121In this series, only a finite number of terms need to be included at a given order in the EFT expansion, because the coefficient of the $U^{k}$ term is already of order ${\cal O}(\mu^{-2k})$.
Indeed, it is not an accident that the first line in (4.45) resembles the four-point function in the EFT, but rather a natural expectation from its analytic structure. Specifically, the corresponding contribution to the four-point function, contained in $\sum_{\pm}x_{0}^{\pm i\omega}F_{\pm 1}^{\text{ansatz}}$, is analytic in the magnitude of the exchanged momentum $k_{3}$. As a result, these parts correspond to the local imprints of the heavy field exchange. [^22] As we elaborated below (3.34), this contribution originates from substituting the time-ordered propagator within $F_{++}$ with the retarded propagator $G_{R}$, thereby capturing all local effects associated with the heavy field exchange. In terms of the four-point function kinematics, this contribution would be analytic in the exchanged momentum $\bm{s}$. However, after sending $\bm{k}_{4}\to 0$, analyticity remains only in $\sqrt{\bm{s}.\bm{s}}=k_{3}$. By contrast, the remaining terms in $F^{\text{ansatz}}$ and the homogeneous part $F^{\text{hom}}_{\pm}(U)$ display non-analyticities proportional to $s^{\pm i\mu}$ and $s^{\pm i\mu\pm i\omega}$, respectively. In consequence, they encode the non-local imprints of the heavy field, which are invisible at any finite order in the EFT expansion. However, as explained in Section 3.3, these non-analytic terms are uniquely fixed by the factorisation property (3.32) in terms of the three-point building block $f$.
@@
which, unlike the previous two cases, is entirely obscured within the structure of $F^{\text{ansatz}}$ in (4.45). In particular, factorisation requires all power laws of the form $u^{2k+1}u^{\frac{1}{2}\pm i\mu\pm i\omega}$ to cancel from $F^{\text{ansatz}}$, allowing only those with even powers to exist. This becomes evident at the level of the final answer only by Taylor-expanding the generalised hypergeometric functions around $U=0$, and even then, by combining terms at different orders in $n$. It can be shown that the remaining terms match order by order in $u$ on both sides of (4.63).
@@
Let us compare our cosmological collider signals with those predicted in related previous works[52, 55, 92, 28, 61, 60]. These studies focus on diagrams with oscillatory cubic vertices and fixed-mass intermediate lines. In particular, it has been shown that vertex oscillations—unlike our mass modulations at order $g^{2}$—can boost the scale-invariant part of the cosmological collider signal [52, 92, 28]. Meanwhile, the dominant scale-dependent resonant collider signal in these studies takes similar forms to the first and/or the second resonant contribution in (4.65), depending on the specifics of the vertex couplings [^23] We thank Xingang Chen for related discussions on the classical cosmological collider signal (see [55] and references therein).. Away from the ultra-squeezed limit, the waveform of our cosmological collider signal departs from those in previous works (with vertex oscillations), as new distinctive patterns of oscillations and scaling behaviours emerge towards the equilateral limit, see e.g. Figure 7.
@@ 4.3 Observational signatures
As was advocated in Section 2, the single-exchange diagrams of the two- and three-point functions of $\pi$ can all be obtained by acting with appropriate weight-shifting operators on the corresponding seed exchange diagram. In this section, for concreteness, we derive these weight-shifting operators assuming $\pi^{\prime}\sigma$ and $\pi^{\prime 2}\sigma$ as the dominant interacting terms in the Lagrangian. Using these vertices, we will go ahead and evaluate the bispectrum using our analytical results for the seed exchange diagram, and for sample values of the parameters $\{g,m_{0},\omega\}$. Finally, we provide explicit formulas for the resonant cosmological collider signal in the squeezed limit of the bispectrum in our setup.
@@
Consider the single-exchange diagram depicted in Figure 2 involving three external massless legs, assuming $\lambda_{R}(\eta)\dot{\pi}_{c}\sigma$ and $\lambda_{L}(\eta)\dot{\pi}_{c}^{2}\sigma$ as its right/left vertices, where $\pi_{c}$ is the canonically normalised Goldstone. For now, we leave their time-dependence generic, but later focus only on constant vertices. The corresponding Schwinger-Keldysh integrand can be mapped onto that of a seed four-point function with the simpler vertices $\lambda_{L,R}(\eta)\varphi^{2}\sigma$. To make this concrete, let us look at the integrand of the bispectrum diagram,
@@
Similarly, the integrands associated with other exchange diagrams (with different vertex structures such as with $\pi\dot{\pi}\sigma,(\partial_{i}\pi)^{2}\sigma,\text{etc}$.) can be mapped onto $\mathbb{F}$, leading to analogous relations between the corresponding power spectrum or the bispectrum and $F$.
@@
Furthermore, although having been ignored in our tree-level computation, there are additional mixings [^25] Such mixings contribute for instance to the double-massive exchange diagram with the $\pi\sigma^{2}$ type vertex, or one-loop contributions to the power spectrum and higher point functions, as studied in [48]. between $\pi$ and $\sigma$ induced by the ever-present operator,
@@
which should be under perturbative control for our setup to be consistent. The conservative  [^26] This is a naive upper bound on the scale of new physics merely on dimensional grounds. Indeed, a more rigorous derivation of the strong coupling scale, albeit in the single-field setup, has shown a higher cut off[112]. While it is beyond the scope of this work, it would be interesting to revisit the strong coupling scale based on $n\to m$ amplitudes in our setup, along the lines of [105]. cutoff associated with this operator is roughly $f=\dot{\phi}/\omega$. Therefore, we need
@@
again at an order of magnitude level. From (4.77) and the inequality above, we get that the condition (4.76)—which is essential for keeping the linear mixing perturbative—is the most stringent lower bound on $\Lambda^{\prime}$. By contrast, the coefficient of the $\dot{\pi}^{2}\sigma$ operator, being independent of the linear mixing, is only constrained by perturbative unitarity, leading to the less stringent bound (4.77) on $\Lambda$. In other words, without an approximate boost symmetry tying their coefficients, the cubic term $\dot{\pi}^{2}\sigma$ consistent with these bounds can be much greater than the $(\partial_{i}\pi)^{2}\sigma$ term, making exchange diagrams involving the latter comparatively negligible. Motivated by this observation, we henceforth only keep the $\dot{\pi}\sigma$ and $\dot{\pi}^{2}\sigma$ vertices in (4.74), while dropping the gradient term $(\partial_{i}\dot{\pi})^{2}\sigma$ for concreteness. We also note that the EFT cubic operator $\pi\pi^{\prime}\sigma$ in (2.6) need not be included, as the vertices (2.6) under consideration are time-independent. Around generic rapidly oscillating backgrounds, however, this term could dominate the cubic action, as recently noted in [92]. Both terms can nevertheless be included in the computation by appropriately adjusting the weight-shifting operator that follows.
@@
These resonant features cannot be arbitrarily large, as the correction to the power spectrum is strongly constrained by observations. To obtain the two-point function, we simply take the limit of the three-point seed function $F({k_{12},s})$ as $k_{3}\to 1$, following the same procedure used in deriving the two-point contact one $f_{\pm 1}(1)$, whose details are provided in the Appendix B. Unlike the case of $f_{\pm 1}(1)$, which can be resummed and expressed compactly in terms of one generalised hypergeometric function of higher weight (4.53), $F(k,k)$ still retains one layer of summation [^28] We find that this series converges rather slowly when the oscillation frequency $\omega$ far exceeds the mass scale $\mu$, as the expression involves extensive products of Gamma and hypergeometric functions. It becomes increasingly difficult to evaluate for larger parameters. We leave the systematic refinement of the series and a search for the closed-form expression through resummation for the future work.. Then, after applying the relevant relation (4.81), we can finally estimate the power spectrum as
$\mathcal{O}(g^{2}):\quad\frac{\Delta P(k)}{P(k)}\sim\mathcal{O}(0.1)\times r_{1}\,g^{2}\,$, (4.94)

### v2

@@ 3.2 Monochromatic masses
around slow-roll backgrounds. Generically, similar oscillations would also appear in the cubic couplings, but to keep the problem’s complexity under control, we take the vertices to be time-independent, setting $\Omega_{L}=\Omega_{R}=0$. Nevertheless, monochromatic vertices could be incorporated into our proposed formulation after minor modifications.
@@
This object is proportional to the cubic wavefunction coefficient, $\psi_{3}=f(k_{12},s)/\sigma_{-}(s,\eta_{\rm end})$ and satisfies the same IDE as (3.18). While we will slightly misuse the terminology and refer to $f$ as the three-point function/building block, the precise relation to the actual three-point function is given by $\langle\phi(\bm{k}_{1})\phi(\bm{k}_{2})\sigma(-\bm{s},\eta_{0})\rangle^{\prime}=\frac{\eta_{\rm end}^{2}}{2k_{1}2k_{2}}\text{Re}\{f(k_{12},s)\sigma_{-}(s,\eta_{0})\}$.
@@ 3.3 Symmetries, microcausality and analyticity
The transformation rule for e.g. $F_{++}$ can be made manifest by writing {keyeqn}
@@
As a side note, the full four-point $F(k_{12},k_{34},s)$ of course admits the same decomposition as (3.25). However, to solve their integro-differential equations, it will be more useful to decompose each Schwinger-Keldysh component separately, as done above.
@@
Inserting the above mode expansions into the respective IDEs does not yield any major simplification, other than producing a recursive set of IDEs for $F_{l}$ and $f_{l}$. Despite this recursive nature, the decomposition becomes especially convenient in perturbation theory, where—up to a fixed order in $g^{2}$—it is sufficient to include only a finite number of harmonics in $l$. To make this simplification more manifest, one can associate the ${\cal O}(g^{2l})$ contribution with a perturbative graph, as in Figure 3, consisting of $l$ mass insertion vertices of the form $\Delta m^{2}(t)\,\sigma^{2}$. This structure implies the following scaling behaviour:
$F_{l}(u)\sim{\cal O}(g^{2|l|})\,,\qquad f_{l}(u)\sim{\cal O}(g^{2|l|})\,$, (3.28)
which ensures that, to describe the exchange diagram $F_{++}$ and its three-point building block $f$ at order ${\cal O}(g^{2})$, we only need to retain the three leading harmonics, $l=0,\pm 1$.
@@
where we have assumed $k_{12},k_{34}$ and $s$ all to be positive. Quite remarkably, the above factorisation property is non-perturbative in $g^{2}$ because it solely relies on the commutation of the heavy field outside the lightcone, independently of its time-dependent mass. In other words, the detailed knowledge of the mode function is not necessary to arrive at (3.33).
However, the derivation of (3.33) relied on the critical assumption that the time integration itself does not introduce additional singularities in $\bm{s}$. One can verify this assumption within a sufficiently small radius around $\bm{s}=0$, where the bulk-bulk propagator $G_{++}$ can be expanded into a power series in $s^{2}$. This expansion typically includes both integer and fractional powers of $s^{2}$. After substituting the propagator with this Taylor expansion, the corresponding Schwinger-Keldysh integral, thanks to the $i\epsilon$ prescription, will still converge for $k_{12},k_{34}>0$, without introducing new singularities near $\bm{s}=0$. As a result, the final answer $F_{++}$ remains analytic around $\bm{s}=0$ except for branch-point singularities descending exclusively from the Wightman propagator at fixed conformal times $\eta,\eta^{\prime}$. Notably, away from the origin $\bm{s}=0$, additional partial energy singularities emerge from time integration, as will be discussed shortly.
We note in passing that (3.33) effectively generalises the cutting rule for non-local cosmological collider signals [70, 19, 72] to scale-breaking cases.
For computing the bispectrum and the power spectrum in our setup, it is sufficient to solve the seed exchange diagram $F_{++}$ only for soft configurations with $\bm{k}_{4}=0$ (hence $k_{34}=k_{3}$ and $s=k_{3}$). So it is convenient to have a factorisation theorem around $\bm{k}_{3}=0$ directly for $F(k_{12},k_{3},k_{3})$. In fact, $F(k_{12},k_{3},k_{3})$ turns out to factorise in this limit in almost the same way as (3.33): {keyeqn}
@@
except that the second contribution is now analytic in $k_{3}$, not necessarily in $k_{3}^{2}$, as (3.33) would have naively suggested. Indeed, since this non-factorisable contribution originates from replacing the Feynman propagator $G_{++}$ with the retarded one $G_{R}$ in $F_{++}$, it inherits odd powers of $k_{3}$ from the right-vertex bulk-boundary propagator($\propto e^{ik_{34}\eta^{\prime}}=\,e^{ik_{3}\,\eta^{\prime}}$). In more detail, the time ordering $\eta^{\prime}>\eta$ enforced by the retarded propagator $G_{R}$ guarantees the convergence of the integral over $\eta^{\prime}$, even after substituting $e^{ik_{3}\,\eta^{\prime}}$ with its Taylor series within the Schwinger-Keldysh integrand. From this substitution, additional odd powers of $k_{3}=\sqrt{k_{3}^{2}}$ appear, multiplying even powers of $k_{3}$ already present in $G_{R}$. This explains the difference between (3.33) and (3.34). Fortunately, odd powers of $k_{3}$ are the only type of non-analyticities that taking the soft limit $\bm{k}_{4}\to 0$ in $F_{++}$ can ever introduce; fractional powers of $k_{3}$ should still factorise in $F_{++}$ according to (3.34). This factorisation will serve as a powerful bootstrap condition for solving the IDEs.
@@ 4.1 Three-point contact diagram
and
@@
Notice that $\hat{\Delta}_{u}$ here differs from its conventional definition in (1.1) by the inclusion of the mass term.
The scale-invariant sector $f_{0}$, which is governed by (4.3), receives no contribution [^8] Note that, despite breaking the dilation symmetry, mass oscillations starting at order $g^{4}$ do contribute to the scale-invariant part of the three-point function because with an even number of mass insertions in the perturbative expansion of $f$, the comoving scale $s$ can partially cancel out between the oscillatory vertices. This will lead to a finite contribution to the scale-invariant component $f_{0}(u)$ at each even order in $g^{2}$ which can be partially understood as mass corrections (see Appendix C). at this order from mass oscillations. Therefore, it is equal to the corresponding three-point function in de Sitter (see e.g. [47, 1]),
@@
which is carried over to the prefactor of the soft term $u^{1/2+i\mu}$ in (4.23) by integration over time. As a result,
@@
is also exponentially small. In fact, both coefficients in (4.23) are exponentially suppressed but for different reasons: the $u^{1/2+i\mu}$ term is small because it is proportional to $\beta_{0}$, while the integration over the late time oscillations of the heavy field suppresses the coefficient of the $u^{1/2-i\mu}$ term.
@@
Let us focus on the mass range $H\ll m_{0}\lesssim\omega$. In this parametric regime, the Bogoliubov coefficient $\Delta\beta$ need not be exponentially small since particles even heavier than $H$ can still be produced by drawing energy from mass oscillations (or equivalently, from the inflaton’s kinetic energy). Indeed, the explicit formula for $\chi^{-}_{1}$ shows a power-law enhancement in $\Delta\beta_{-}$:
@@ 4.2 Three-point exchange diagram
    The particular solution (4.45) represent two qualitatively different contributions to the exchange diagram, $F_{++}=x_{0}^{\pm i\omega}F_{\pm}^{\text{ansatz}}$, as a function of the intermediate momentum $k_{3}$. The first line’s contribution is analytic around $k_{3}=0$ since factors of $k_{3}^{\pm i\omega}$ are precisely cancelled between the prefactor $x_{0}^{\pm i\omega}$ and $F^{\text{ansatz}}_{\pm 1}$. However, the second line in (4.45) induces a branch point at $k_{3}=0$ in the final four-point function. We shall soon come to the physical interpretation of each analytic behaviour below. These coefficients have a pole at $\omega=2\mu$, indicating a breakdown of perturbative calculation. As will be shown in Section 5.1, this singluar behaviour is in fact a manifestation of resonance that is a non-perturbative effect beyond the reach of perturbative solutions. Further discussion of the radius of convergence of the perturbative expansion is provided in Appendix C.
@@
    in which $_{p}\tilde{\rm{F}}_{q}$ is the regularised hypergeometric function defined in (1.5). Let us stress that while it was sufficient to take the leading soft limit $k_{3}\to 0$ to fix the free coefficients $\xi_{1,2}$, the factorisation in (3.34) holds to all orders in $k_{3}$. We shall elaborate on this point below. Additional intermediate steps leading to the above results are spelled out in Appendix B.
@@
while $\mathcal{A}^{\pm}_{n>0}$ are of order $1/\mu^{4}$ or higher. By plugging the asymptotic form of $\mathcal{A}^{\pm}_{0}$ into (4.45) and neglecting the remaining terms, we recover the leading order EFT term $F^{\text{EFT}}_{\pm 1}$. Similarly, higher order EFT contributions could be isolated in (4.45) by keeping more powers of $1/\mu^{2}$ in the prefactors $\mathcal{A}^{\pm}_{n}$, and by expanding $_{3}F_{2}$ near $U=0$. In this series, only a finite number of terms need to be included at a given order in the EFT expansion, because the coefficient of the $U^{k}$ term is already of order ${\cal O}(\mu^{-2k})$.
Indeed, it is not an accident that the first line in (4.45) resembles the four-point function in the EFT, but rather a natural expectation from its analytic structure. Specifically, the corresponding contribution to the four-point function, contained in $\sum_{\pm}x_{0}^{\pm i\omega}F_{\pm 1}^{\text{ansatz}}$, is analytic in the magnitude of the exchanged momentum $k_{3}$. As a result, these parts correspond to the local imprints of the heavy field exchange. As we elaborated below (3.34), this contribution originates from substituting the time-ordered propagator within $F_{++}$ with the retarded propagator $G_{R}$, thereby capturing all local effects associated with the heavy field exchange. In terms of the four-point function kinematics, this contribution would be analytic in the exchanged momentum $\bm{s}$. However, after sending $\bm{k}_{4}\to 0$, analyticity remains only in $\sqrt{\bm{s}.\bm{s}}=k_{3}$. By contrast, the remaining terms in $F^{\text{ansatz}}$ and the homogeneous part $F^{\text{hom}}_{\pm}(U)$ display non-analyticities proportional to $s^{\pm i\mu}$ and $s^{\pm i\mu\pm i\omega}$, respectively. These are, correspondingly, the non-local imprints of the exchange of the heavy field, which remain invisible at any finite order in the EFT expansion. However, as explained in Section 3.3, these non-analytic terms are uniquely fixed in terms of the three-point building block $f$, thanks to the factorisation property (3.32).
@@
which, unlike the previous two cases, is entirely obscured within the structure of $F^{\text{ansatz}}$ in (4.45). In particular, factorisation requires all power laws of the form $u^{2k+1}u^{\frac{1}{2}\pm i\mu\pm i\omega}$ to cancel from $F^{\text{ansatz}}$, leaving only terms with even powers. This cancellation becomes evident at the level of the final answer after Taylor-expanding the generalised hypergeometric functions around $U=0$, and even then, only by combining terms at different orders in $n$. It can be shown that the remaining terms match order by order in $u$ on both sides of (4.63).
@@
Let us compare our cosmological collider signals with those predicted in related previous works[53, 56, 93, 28, 62, 61]. These studies focus on diagrams with oscillatory cubic vertices and fixed-mass intermediate lines. In particular, it has been shown that vertex oscillations—unlike our mass modulations at order $g^{2}$—can boost the scale-invariant part of the cosmological collider signal [53, 93, 28]. Meanwhile, the dominant scale-dependent resonant collider signal in these studies takes similar forms to the first and/or the second resonant contribution in (4.65), depending on the specifics of the vertex couplings [^10] We thank Xingang Chen for related discussions on the classical cosmological collider signal (see [56] and references therein).. Away from the ultra-squeezed limit, the waveform of our cosmological collider signal departs from those presented in previous works (with vertex oscillations), as new distinctive patterns of oscillations and scaling behaviours emerge towards the equilateral limit, see e.g. Figure 7.
@@ 4.3 Observational signatures
As was advocated in Section 2, the single-exchange diagrams of the two- and three-point functions of $\pi$ can all be obtained by acting with appropriate weight-shifting operators on the corresponding seed exchange diagram. In this section, for concreteness, we derive these weight-shifting operators assuming $\pi^{\prime}\sigma$ and $\pi^{\prime 2}\sigma$ are the dominant interacting terms in the Lagrangian. Using these vertices, we will go ahead and evaluate the bispectrum using our analytical results for the seed exchange diagram, and for sample values of the parameters $\{g,m_{0},\omega\}$. Finally, we provide explicit formulas for the resonant cosmological collider signal in the squeezed limit of the bispectrum in our setup.
@@
Consider the single-exchange diagram depicted in Figure 2 involving three external massless legs, assuming $\lambda_{R}(\eta)\dot{\pi}_{c}\sigma$ and $\lambda_{L}(\eta)\dot{\pi}_{c}^{2}\sigma$ as its right/left vertices, where $\pi_{c}$ is the canonically normalised Goldstone. For now, we leave their time-dependence generic, but later restrict to constant vertices. The corresponding Schwinger-Keldysh integrand can be mapped onto that of a seed four-point function with the simpler vertices $\lambda_{L,R}(\eta)\varphi^{2}\sigma$. To make this point concrete, let us look at the integrand of the bispectrum diagram,
@@
Similarly, the integrands associated with other exchange diagrams (with different vertex structures such as $\pi\dot{\pi}\sigma,(\partial_{i}\pi)^{2}\sigma,\text{etc}$.) can be mapped onto $\mathbb{F}$, leading to analogous relations between the corresponding power spectrum or the bispectrum and $F$.
@@
Furthermore, although having been ignored in our tree-level computation, there are additional mixings between $\pi$ and $\sigma$ induced by the ever-present operator,
@@
which should be under perturbative control for our setup to be consistent. Such mixings contribute for instance to the double-massive exchange diagram with the $\pi\sigma^{2}$ type vertex, or one-loop contributions to the power spectrum and higher point functions, as studied in [49].
The conservative value for the cutoff associated with this operator is roughly $f=\dot{\phi}/\omega$, which is a naive upper bound on the scale of new physics merely on dimensional grounds. A more rigorous derivation in the single-field setup has shown a higher cutoff [113]. While it is beyond the scope of this work, it would be interesting to revisit this strong coupling scale using $n\to m$ amplitudes in our setup, along the lines of [106]. In the following, we adopt this naive cutoff, which requires us to impose
@@
at an order of magnitude level.
From (4.77) and the inequality above, we get that the condition (4.76)—which is essential for keeping the linear mixing perturbative—is the most stringent lower bound on $\Lambda^{\prime}$. By contrast, the coefficient of the $\dot{\pi}^{2}\sigma$ operator, being independent of the linear mixing, is only constrained by perturbative unitarity, leading to the less stringent bound (4.77) on $\Lambda$. In other words, without an approximate boost symmetry tying their coefficients, the cubic term $\dot{\pi}^{2}\sigma$ consistent with these bounds can be much greater than the $(\partial_{i}\pi)^{2}\sigma$ term, making exchange diagrams involving the latter comparatively negligible. Motivated by this observation, we henceforth only keep the $\dot{\pi}\sigma$ and $\dot{\pi}^{2}\sigma$ vertices in (4.74), while dropping the gradient term $(\partial_{i}\dot{\pi})^{2}\sigma$ for concreteness. We also note that the EFT cubic operator $\pi\pi^{\prime}\sigma$ in (2.6) need not be included, as the vertices (2.6) under consideration are time-independent. However, around generic rapidly oscillating backgrounds, this term could dominate the cubic action, as recently noted in [93]. Both terms can nevertheless be included in our computation by appropriately adjusting the weight-shifting operator discussed below.
@@
These resonant features cannot be arbitrarily large, as the correction to the power spectrum is strongly constrained by observations. To obtain the two-point function, we simply take the limit of the three-point seed function $F({k_{12},s})$ as $k_{3}\to 1$, following the same procedure used in deriving the two-point contact one $f_{\pm 1}(1)$, whose details are provided in the Appendix B. Unlike the case of $f_{\pm 1}(1)$, which can be resummed and expressed compactly in terms of one generalised hypergeometric function of higher weight (4.53), $F(k,k)$ still retains one layer of summation. We find that this series converges rather slowly when the oscillation frequency $\omega$ far exceeds the mass scale $\mu$, as the expression involves extensive products of Gamma and hypergeometric functions. It becomes increasingly difficult to evaluate for larger parameters. We leave the systematic refinement of the series and a search for the closed-form expression through resummation for future work.
After applying the relevant relation (4.81), we can finally estimate the power spectrum as
$\mathcal{O}(g^{2}):\quad\frac{\Delta P(k)}{P(k)}\sim\mathcal{O}(0.1)\times r_{1}\,.g^{2}\,$, (4.94)